A firm producing five units of output has an average total cost of R100 and has to pay
R150 to its fixed factors of production. The average variable cost is …

1. R30
2. R70
3. R100
4. R150
1
Expert's answer
2019-02-25T14:37:08-0500

Fixed cost per unit of output:150/5 = R 30

Variable cost per unit of output: ttl cost- fixed cost =100-30 = R 70